# thanks: Corresponding author

### v1

@@ III Numerical results and discussions
$\Xi_{cc}^{*}(\frac{3}{2}^{+})(1S)$ 3.506.003.50-6.00 $3.93^{2}-4.30^{2}$ $3.77_{-0.02}^{+0.03}$ $0.17_{-0.01}^{+0.02}$
$\Xi_{cc}^{*}$ $\Xi_{cc}^{*}(\frac{3}{2}^{-})(1P)$ 3.506.003.50-6.00 $4.43^{2}-4.78^{2}$ $3.95_{-0.04}^{+0.03}$ $0.19\pm 0.02$
$\Xi_{cc}^{*}(\frac{3}{2}^{+})(2S)$ 3.506.003.50-6.00 $4.91^{2}-5.29^{2}$ $4.10\pm 0.04$ $0.21\pm 0.02$
$\Xi_{bc}^{*}(\frac{3}{2}^{+})(1S)$ 6.009.006.00-9.00 $6.96^{2}-7.25^{2}$ $6.97\pm 0.06$ $0.31\pm 0.03$
$\Xi_{bc}^{*}$ $\Xi_{bc}^{*}(\frac{3}{2}^{-})(1P)$ 6.009.006.00-9.00 $7.46^{2}-7.75^{2}$ $7.19\pm 0.07$ $0.35\pm 0.03$
$\Xi_{bc}^{*}(\frac{3}{2}^{+})(2S)$ 6.009.006.00-9.00 $7.96^{2}-8.25^{2}$ $7.38\pm 0.07$ $0.37_{-0.03}^{+0.04}$
$\Xi_{bb}^{*}(\frac{3}{2}^{+})(1S)$ 10.0015.0010.00-15.00 $10.68^{2}-10.96^{2}$ $10.18\pm 0.10$ $0.49\pm 0.04$
$\Xi_{bb}^{*}$ $\Xi_{bb}^{*}(\frac{3}{2}^{-})(1P)$ 10.0015.0010.00-15.00 $11.18^{2}-11.46^{2}$ $10.39\pm 0.12$ $0.66_{-0.04}^{+0.06}$
$\Xi_{bb}^{*}(\frac{3}{2}^{+})(2S)$ 10.0015.0010.00-15.00 $11.68^{2}-11.96^{2}$ $10.60\pm 0.12$ $0.75\pm 0.06$
$\Omega_{cc}^{*}(\frac{3}{2}^{+})(1S)$ 3.506.003.50-6.00 $3.94^{2}-4.31^{2}$ $3.79\pm 0.03$ $0.18\pm 0.02$
$\Omega_{cc}^{*}$ $\Omega_{cc}^{*}(\frac{3}{2}^{-})(1P)$ 3.506.003.50-6.00 $4.44^{2}-4.79^{2}$ $3.99_{-0.04}^{+0.03}$ $0.21_{-0.01}^{+0.02}$
$\Omega_{cc}^{*}(\frac{3}{2}^{+})(2S)$ 3.506.003.50-6.00 $4.92^{2}-5.30^{2}$ $4.14_{-0.04}^{+0.03}$ $0.22\pm 0.02$
$\Omega_{bc}^{*}(\frac{3}{2}^{+})(1S)$ 6.009.006.00-9.00 $6.96^{2}-7.25^{2}$ $7.04_{-0.05}^{+0.06}$ $0.33\pm 0.03$
$\Omega_{bc}^{*}$ $\Omega_{bc}^{*}(\frac{3}{2}^{-})(1P)$ 6.009.006.00-9.00 $7.46^{2}-7.75^{2}$ $7.26\pm 0.07$ $0.37_{-0.03}^{+0.05}$
$\Omega_{bc}^{*}(\frac{3}{2}^{+})(2S)$ 6.009.006.00-9.00 $7.96^{2}-8.25^{2}$ $7.45\pm 0.07$ $0.39\pm 0.04$
$\Omega_{bb}^{*}(\frac{3}{2}^{+})(1S)$ 10.0015.0010.00-15.00 $10.68^{2}-10.96^{2}$ $10.20\pm 0.10$ $0.50\pm 0.04$
$\Omega_{bb}^{*}$ $\Omega_{bb}^{*}(\frac{3}{2}^{-})(1P)$ 10.0015.0010.00-15.00 $11.18^{2}-11.46^{2}$ $10.44\pm 0.12$ $0.67\pm 0.05$
$\Omega_{bb}^{*}(\frac{3}{2}^{+})(2S)$ 10.0015.0010.00-15.00 $11.68^{2}-11.96^{2}$ $10.66\pm 0.12$ $0.77\pm 0.06$
@@
As presented in table 3, the masses of the first orbital $1P$ and radial $2S$ excited states of $\Xi_{cc}^{*}$ exceed that of the ground state by approximately 0.18 and 0.33 $\mathrm{GeV}$, respectively. Corresponding mass differences for the other baryons are: $\Xi_{bc}^{*}$ (0.22 and 0.41 $\mathrm{GeV}$), $\Xi_{bb}^{*}$ (0.21 and 0.42 $\mathrm{GeV}$), $\Omega_{cc}^{*}$ (0.20 and 0.35 $\mathrm{GeV}$), $\Omega_{bc}^{*}$ (0.22 and 0.41 $\mathrm{GeV}$), and $\Omega_{bb}^{*}$ (0.22 and 0.46 $\mathrm{GeV}$).
Table 4 provides a comprehensive comparison of the predicted masses for the spin-$\frac{3}{2}$ doubly heavy baryons in the $1S$ state, derived using various theoretical frameworks. Our analysis incorporates nonperturbative operators with mass dimensions extended up to 10, aiming to enhance precision and minimize uncertainties. This approach advances previous studies, such as the QCD sum rule analysis in Aliev:2012iv , which considered operators only up to dimension 5. As can be seen in table 4, our predictions are in good agreement with the Lattice QCD results Brown:2014ena . The mass estimates obtained here show consistency with results from alternative techniques, including the relativistic quark model Ebert:2002ig , QCD sum rule method Aliev:2012iv , and constituent quark models Yoshida:2015tia , within the stated error margins for several channels.
@@
$\Xi_{cc}^{*}$ $3.77_{-0.02}^{+0.03}$ $3.69\pm 0.16$ $3.692\pm 0.028\pm 0.021$ $3.727$ $3.90\pm 0.10$ $3.754$ $3.61\pm 0.18$
$\Xi_{bc}^{*}$ $6.97\pm 0.06$ $7.25\pm 0.20$ $5\pm 0.036\pm 0.028 6.986.98$ $8.00\pm 0.26$ - -
$\Xi_{bb}^{*}$ $10.18\pm 0.10$ $10.4\pm 1.0$ $10.187\pm 0.030\pm 0.024$ $10.237$ $10.35\pm 0.08$ $10.339$ $10.22\pm 0.15$
$\Omega_{cc}^{*}$ $3.79\pm 0.03$ $3.78\pm 0.16$ $3.822\pm 0.030\pm 0.024$ $3.872$ $1\pm 0.06 3.883$3.883 $3.76\pm 0.17$
$\Omega_{bc}^{*}$ $7.04_{-0.05}^{+0.06}$ $7.3\pm 0.2$ $7.059\pm 0.028\pm 0.021$ $7.13$ $7.54\pm 0.08$ - -
$\Omega_{bb}^{*}$ $10.20\pm 0.10$ $10.5\pm 0.2$ $10.308\pm 0.027\pm 0.021$ $10.389$ $10.28\pm 00.05$ $10.467$ $10.38\pm 0.14$
@@
As previously mentioned, the mass spectra of the spin-$\frac{3}{2}$ doubly heavy baryons in the $1P$ and $2S$ states have been explored using various theoretical approaches. Our study provides mass predictions for these excited states with improved precision. The results displayed in tables 5 and 6 demonstrate that, for most channels, our findings align well with those obtained from different frameworks, including the hypercentral constituent quark model Shah:2016vmd ; Shah:2017liu , constituent quark models Yoshida:2015tia , and the QCD sum rule method Alrebdi:2022lat ; Wang:2010it , all within the quoted uncertainties.
@@
$\Xi_{cc}^{*}$ $3.95_{-0.04}^{+0.03}$ $3.85$ - $3.80\pm 0.18$ $3.949$ $5.15\pm 0.05$
$\Xi_{bc}^{*}$ $7.19\pm 0.07$ $7.15$ - - - $7.95\pm 0.05$
$\Xi_{bb}^{*}$ $10.39\pm 0.12$ $10.51$ - $10.43\pm 0.15$ $10.476$ $10.95\pm 0.025$
$\Omega_{cc}^{*}$ $3.99_{-0.04}^{+0.03}$ - $3.97$ $3.96\pm 0.16$ $4.086$ $4.95\pm 0.05$
$\Omega_{bc}^{*}$ $7.26\pm 0.07$ - $7.37$ - - $7.85\pm 0.05$
$\Omega_{bb}^{*}$ $10.44\pm 0.12$ - $10.64$ $10.57\pm 0.15$ $10.608$ $10.85\pm 0.025$
@@
$\Xi_{cc}^{*}$ $4.10\pm 0.04$ $3.99$ - $4.95\pm 0.05$
$\Xi_{bc}^{*}$ $7.38\pm 0.07$ $7.27$ - $7.91\pm 0.05$
$\Xi_{bb}^{*}$ $10.60\pm 0.12$ $10.62$ - $10.92\pm 0.025$
$\Omega_{cc}^{*}$ $4.14_{-0.04}^{+0.03}$ - $4.10$ $5.00\pm 0.05$
$\Omega_{bc}^{*}$ $7.45\pm 0.07$ - $7.48$ $7.90\pm 0.05$
$\Omega_{bb}^{*}$ $10.66\pm 0.12$ - $10.74$ $10.90\pm 0.025$
@@
This study also analyzes the residues associated with the ground state, the first orbital $1P$ excitation, and the first radial $2S$ excitation of doubly heavy baryons possessing spin-$\frac{3}{2}$, with corresponding data presented in tables 78, and 9. Although earlier works using QCD sum rules have investigated the residues in the ground state Aliev:2012iv ; Wang:2010hs ; Bagan:1992za , as well as the $1P$ Alrebdi:2022lat ; Wang:2010it and $2S$ excited states Alrebdi:2022lat , our current findings offer improved accuracy for residue calculations. Our calculations yield enhanced precision for the residue values compared to prior predictions, owing to the inclusion of nonperturbative operators up to dimension ten in the OPE. Although mass predictions across various approaches generally exhibit strong agreement (within reported uncertainties, with minor channel specific differences), we observe considerable variations in the reported residue values for $1S$, $1P$, and $2S$ states among different studies, despite some channel specific consistencies. The inherently larger uncertainties associated with the presented residues, relative to their corresponding masses, are a consequence of their derivation: masses are obtained from a ratio of sum rules, which naturally mitigates uncertainties, whereas residues are directly extracted from a single sum rule. These newly presented residue values offer crucial input for future investigations into the decay channels of the doubly heavy baryons.
Baryon Our work Aliev:2012iv Wang:2010hs Bagan:1992za
$\Xi_{cc}^{*}$ $0.17_{-0.01}^{+0.02}$ $12\pm 0.01 0.0$$70\pm 0.017 0.0$$071\pm 0.017$
$\Xi_{bc}^{*}$ $0.31\pm 0.03$ $0.15\pm 0.01$ - -
$\Xi_{bb}^{*}$ $0.49\pm 0.04$ $0.22\pm 0.03$ $0.161\pm 0.041$ $0.111\pm 0.040$
$\Omega_{cc}^{*}$ $0.18\pm 0.02$ $14\pm 0.02 0.0$$085\pm 0.019 -$
$\Omega_{bc}^{*}$ $0.33\pm 0.03$ $0.18\pm 0.02$ - -
$\Omega_{bb}^{*}$ $0.50\pm 0.04$ $0.25\pm 0.03$ $0.199\pm 0.048$ -
@@
$\Xi_{cc}^{*}$ $19\pm 0.02 0.0$$095\pm 0.020$ $0.185\pm 0.025$
$\Xi_{bc}^{*}$ $0.35\pm 0.03$ - $0.480\pm 0.025$
$\Xi_{bb}^{*}$ $0.66_{-0.04}^{+0.06}$ $0.227\pm 0.052$ $1.060\pm 0.05$
$\Omega_{cc}^{*}$ $0.21_{-0.01}^{+0.02}$ $0.116\pm 0.022$ $0.165\pm 0.025$
$\Omega_{bc}^{*}$ $0.37_{-0.03}^{+0.05}$ - $0.480\pm 0.025$
$\Omega_{bb}^{*}$ $0.67\pm 0.05$ $0.275\pm 0.059$ $1.075\pm 0.05$
@@
$\Xi_{cc}^{*}$ $0.21\pm 0.02$ $0.175\pm 0.025$
$\Xi_{bc}^{*}$ $0.37_{-0.03}^{+0.04}$ $0.475\pm 0.025$
$\Xi_{bb}^{*}$ $0.75\pm 0.06$ $1.050\pm 0.05$
$\Omega_{cc}^{*}$ $0.22\pm 0.02$ $0.160\pm 0.025$
$\Omega_{bc}^{*}$ $0.39\pm 0.04$ $0.475\pm 0.025$
$\Omega_{bb}^{*}$ $0.77\pm 0.06$ $1.080\pm 0.05$
@@
Accurate spectroscopic parameters, including masses and residues, are essential for investigating the interactions and decays of doubly heavy baryons. Theoretical calculations of these parameters also serve as crucial guides for experimental searches for yet-undiscovered baryon states predicted by quark models. This study focuses on precisely determining the spectroscopic parameters of doubly heavy baryons across the ground state ($1S$), first orbital excitation ($1P$), and the first radial excitation ($2S$). Using the QCD sum rule approach, a reliable and predictive technique based on the QCD Lagrangian, we include nonperturbative operators up to dimension ten to enhance accuracy. Following the determination of auxiliary parameters, we evaluate the mass spectra and residue for these specific baryon states, presenting them in table 3 and comparing our findings with results from other theoretical frameworks. The outcomes of this study are poised to aid experimental groups in their quest to discover the yet-unseen members of the doubly heavy baryon family. Additionally, these predictions offer a valuable basis for comparison with future theoretical work employing different nonperturbative methods. The importance of residue values, critical for computing any physical quantities related to baryon interactions and decays, cannot be overstated. Our high accuracy results provide the necessary input data for comprehensive studies of the decay characteristics of the considered doubly heavy baryons in their ground and excited configurations. These investigations are essential for determining the decay widths and estimating the lifetimes of these baryons.

### v2

@@ III Numerical results and discussions
$\Xi_{cc}^{*}(\frac{3}{2}^{+})(1S)$ 3.506.003.50-6.00 $3.93^{2}-4.30^{2}$ $3.68_{-0.14}^{+0.15}$ $0.15\pm 0.02$
$\Xi_{cc}^{*}$ $\Xi_{cc}^{*}(\frac{3}{2}^{-})(1P)$ 3.506.003.50-6.00 $4.43^{2}-4.78^{2}$ $3.85_{-0.15}^{+0.16}$ $0.17\pm 0.02$
$\Xi_{cc}^{*}(\frac{3}{2}^{+})(2S)$ 3.506.003.50-6.00 $4.91^{2}-5.29^{2}$ $3.99\pm 0.16$ $0.19\pm 0.03$
$\Xi_{bc}^{*}(\frac{3}{2}^{+})(1S)$ 6.009.006.00-9.00 $6.96^{2}-7.25^{2}$ $6.95\pm 0.20$ $0.30\pm 0.03$
$\Xi_{bc}^{*}$ $\Xi_{bc}^{*}(\frac{3}{2}^{-})(1P)$ 6.009.006.00-9.00 $7.46^{2}-7.75^{2}$ $7.16\pm 0.21$ $0.34\pm 0.03$
$\Xi_{bc}^{*}(\frac{3}{2}^{+})(2S)$ 6.009.006.00-9.00 $7.96^{2}-8.25^{2}$ $7.37\pm 0.21$ $0.36_{-0.03}^{+0.04}$
$\Xi_{bb}^{*}(\frac{3}{2}^{+})(1S)$ 10.0015.0010.00-15.00 $10.68^{2}-10.96^{2}$ $10.28\pm 0.30$ $0.50\pm 0.05$
$\Xi_{bb}^{*}$ $\Xi_{bb}^{*}(\frac{3}{2}^{-})(1P)$ 10.0015.0010.00-15.00 $11.18^{2}-11.46^{2}$ $10.48\pm 0.32$ $0.67_{-0.05}^{+0.06}$
$\Xi_{bb}^{*}(\frac{3}{2}^{+})(2S)$ 10.0015.0010.00-15.00 $11.68^{2}-11.96^{2}$ $10.71\pm 0.32$ $0.77\pm 0.07$
$\Omega_{cc}^{*}(\frac{3}{2}^{+})(1S)$ 3.506.003.50-6.00 $3.94^{2}-4.31^{2}$ $3.74\pm 0.15$ $0.16\pm 0.02$
$\Omega_{cc}^{*}$ $\Omega_{cc}^{*}(\frac{3}{2}^{-})(1P)$ 3.506.003.50-6.00 $4.44^{2}-4.79^{2}$ $3.96_{-0.16}^{+0.15}$ $0.20_{-0.02}^{+0.03}$
$\Omega_{cc}^{*}(\frac{3}{2}^{+})(2S)$ 3.506.003.50-6.00 $4.92^{2}-5.30^{2}$ $4.10_{-0.17}^{+0.16}$ $0.21\pm 0.03$
$\Omega_{bc}^{*}(\frac{3}{2}^{+})(1S)$ 6.009.006.00-9.00 $6.96^{2}-7.25^{2}$ $7.01_{-0.19}^{+0.20}$ $0.33\pm 0.04$
$\Omega_{bc}^{*}$ $\Omega_{bc}^{*}(\frac{3}{2}^{-})(1P)$ 6.009.006.00-9.00 $7.46^{2}-7.75^{2}$ $7.21\pm 0.21$ $0.36_{-0.04}^{+0.05}$
$\Omega_{bc}^{*}(\frac{3}{2}^{+})(2S)$ 6.009.006.00-9.00 $7.96^{2}-8.25^{2}$ $7.40\pm 0.21$ $0.38\pm 0.05$
$\Omega_{bb}^{*}(\frac{3}{2}^{+})(1S)$ 10.0015.0010.00-15.00 $10.68^{2}-10.96^{2}$ $10.34\pm 0.30$ $0.53\pm 0.05$
$\Omega_{bb}^{*}$ $\Omega_{bb}^{*}(\frac{3}{2}^{-})(1P)$ 10.0015.0010.00-15.00 $11.18^{2}-11.46^{2}$ $10.60\pm 0.32$ $0.70\pm 0.06$
$\Omega_{bb}^{*}(\frac{3}{2}^{+})(2S)$ 10.0015.0010.00-15.00 $11.68^{2}-11.96^{2}$ $10.82\pm 0.32$ $0.80\pm 0.08$
@@
As presented in table 2, the masses of the first orbital $1P$ and radial $2S$ excited states of $\Xi_{cc}^{*}$ exceed that of the ground state by approximately 0.17 and 0.31 $\mathrm{GeV}$, respectively. Corresponding mass differences for the other baryons are: $\Xi_{bc}^{*}$ (0.21 and 0.42 $\mathrm{GeV}$), $\Xi_{bb}^{*}$ (0.20 and 0.43 $\mathrm{GeV}$), $\Omega_{cc}^{*}$ (0.22 and 0.36 $\mathrm{GeV}$), $\Omega_{bc}^{*}$ (0.20 and 0.39 $\mathrm{GeV}$), and $\Omega_{bb}^{*}$ (0.26 and 0.48 $\mathrm{GeV}$).
Table 3 provides a comprehensive comparison of the predicted masses for the spin-$\frac{3}{2}$ doubly heavy baryons in the $1S$ state, derived using various theoretical frameworks. Our analysis incorporates nonperturbative operators with mass dimensions extended up to 10, aiming to enhance precision and minimize uncertainties. As can be seen in table 3, our predictions are in good agreement with the Lattice QCD results Brown:2014ena . The mass estimates obtained here show consistency with results from alternative techniques, including the relativistic quark model Ebert:2002ig , QCD sum rule method Aliev:2012iv , and constituent quark models Yoshida:2015tia , within the stated error margins for several channels.
@@
$\Xi_{cc}^{*}$ $3.68_{-0.14}^{+0.15}$ $3.69\pm 0.16$ $3.692\pm 0.028\pm 0.021$ $3.727$ $3.90\pm 0.10$ $3.754$ $3.61\pm 0.18$
$\Xi_{bc}^{*}$ $6.95\pm 0.20$ $7.25\pm 0.20$ $5\pm 0.036\pm 0.028 6.986.98$ $8.00\pm 0.26$ - -
$\Xi_{bb}^{*}$ $10.28\pm 0.30$ $10.4\pm 1.0$ $10.187\pm 0.030\pm 0.024$ $10.237$ $10.35\pm 0.08$ $10.339$ $10.22\pm 0.15$
$\Omega_{cc}^{*}$ $3.74\pm 0.15$ $3.78\pm 0.16$ $3.822\pm 0.030\pm 0.024$ $3.872$ $1\pm 0.06 3.883$3.883 $3.76\pm 0.17$
$\Omega_{bc}^{*}$ $7.01_{-0.19}^{+0.20}$ $7.3\pm 0.2$ $7.059\pm 0.028\pm 0.021$ $7.13$ $7.54\pm 0.08$ - -
$\Omega_{bb}^{*}$ $10.34\pm 0.30$ $10.5\pm 0.2$ $10.308\pm 0.027\pm 0.021$ $10.389$ $10.28\pm 00.05$ $10.467$ $10.38\pm 0.14$
@@
As previously mentioned, the mass spectra of the spin-$\frac{3}{2}$ doubly heavy baryons in the $1P$ and $2S$ states have been explored using various theoretical approaches. Our study provides mass predictions for these excited states with improved precision. The results displayed in tables 4 and 5 demonstrate that, for most channels, our findings align well with those obtained from different frameworks, including the hypercentral constituent quark model Shah:2016vmd ; Shah:2017liu , constituent quark models Yoshida:2015tia , and the QCD sum rule method Wang:2010it , all within the quoted uncertainties.
@@
$\Xi_{cc}^{*}$ $_{-0.15}^{+0.16} 3.853.85$ - $3.80\pm 0.18$ $3.949$ $5.15\pm 0.05$
$\Xi_{bc}^{*}$ $7.16\pm 0.21$ $7.15$ - - - $7.95\pm 0.05$
$\Xi_{bb}^{*}$ $10.48\pm 0.32$ $10.51$ - $10.43\pm 0.15$ $10.476$ $10.95\pm 0.025$
$\Omega_{cc}^{*}$ $3.96_{-0.16}^{+0.15}$ - $3.97$ $3.96\pm 0.16$ $4.086$ $4.95\pm 0.05$
$\Omega_{bc}^{*}$ $7.21\pm 0.21$ - $7.37$ - - $7.85\pm 0.05$
$\Omega_{bb}^{*}$ $10.60\pm 0.32$ - $10.64$ $10.57\pm 0.15$ $10.608$ $10.85\pm 0.025$
@@
$\Xi_{cc}^{*}$ $\pm 0.16 3.993.99$ - $4.95\pm 0.05$
$\Xi_{bc}^{*}$ $7.37\pm 0.21$ $7.27$ - $7.91\pm 0.05$
$\Xi_{bb}^{*}$ $10.71\pm 0.32$ $10.62$ - $10.92\pm 0.025$
$\Omega_{cc}^{*}$ $_{-0.17}^{+0.16} - 4.104.10$ $5.00\pm 0.05$
$\Omega_{bc}^{*}$ $7.40\pm 0.21$ - $7.48$ $7.90\pm 0.05$
$\Omega_{bb}^{*}$ $10.82\pm 0.32$ - $10.74$ $10.90\pm 0.025$
@@
This study also analyzes the residues associated with the ground state, the first orbital $1P$ excitation, and the first radial $2S$ excitation of doubly heavy baryons possessing spin-$\frac{3}{2}$, with corresponding data presented in tables 67, and 8. Although earlier works using QCD sum rules have investigated the residues in the ground state Aliev:2012iv ; Wang:2010vn ; Bagan:1992za , as well as the $1P$ Alrebdi:2022lat ; Wang:2010it and $2S$ excited states Alrebdi:2022lat , our current findings offer improved accuracy for residue calculations. Although mass predictions across various approaches generally exhibit strong agreement (within reported uncertainties, with minor channel specific differences), we observe considerable variations in the reported residue values for $1S$, $1P$, and $2S$ states among different studies, despite some channel specific consistencies. The inherently larger uncertainties associated with the presented residues, relative to their corresponding masses, are a consequence of their derivation: masses are obtained from a ratio of sum rules, which naturally mitigates uncertainties, whereas residues are directly extracted from a single sum rule. These newly presented residue values offer crucial input for future investigations into the decay channels of the doubly heavy baryons.
Baryon Our work Aliev:2012iv Bagan:1992za Wang:2010vn
$\Xi_{cc}^{*}$ $0.15\pm 0.02$ $12\pm 0.01 0.0$$71\pm 0.017 0.0$$070\pm 0.017$
$\Xi_{bc}^{*}$ $0.30\pm 0.03$ $0.15\pm 0.01$ - -
$\Xi_{bb}^{*}$ $0.50\pm 0.05$ $0.22\pm 0.03$ $0.111\pm 0.040$ $0.161\pm 0.041$
$\Omega_{cc}^{*}$ $0.16\pm 0.02$ $14\pm 0.02 - 0.0$$085\pm 0.019$
$\Omega_{bc}^{*}$ $0.33\pm 0.04$ $0.18\pm 0.02$ - -
$\Omega_{bb}^{*}$ $0.53\pm 0.05$ $0.25\pm 0.03$ - $0.199\pm 0.048$
@@
$\Xi_{cc}^{*}$ $17\pm 0.02 0.0$$095\pm 0.020$ $0.185\pm 0.025$
$\Xi_{bc}^{*}$ $0.34\pm 0.03$ - $0.480\pm 0.025$
$\Xi_{bb}^{*}$ $0.67_{-0.05}^{+0.06}$ $0.227\pm 0.052$ $1.060\pm 0.05$
$\Omega_{cc}^{*}$ $0.20_{-0.02}^{+0.03}$ $0.116\pm 0.022$ $0.165\pm 0.025$
$\Omega_{bc}^{*}$ $0.36_{-0.04}^{+0.05}$ - $0.480\pm 0.025$
$\Omega_{bb}^{*}$ $0.70\pm 0.06$ $0.275\pm 0.059$ $1.075\pm 0.05$
@@
$\Xi_{cc}^{*}$ $0.19\pm 0.03$ $0.175\pm 0.025$
$\Xi_{bc}^{*}$ $0.36_{-0.03}^{+0.04}$ $0.475\pm 0.025$
$\Xi_{bb}^{*}$ $0.77\pm 0.07$ $1.050\pm 0.05$
$\Omega_{cc}^{*}$ $0.21\pm 0.03$ $0.160\pm 0.025$
$\Omega_{bc}^{*}$ $0.38\pm 0.05$ $0.475\pm 0.025$
$\Omega_{bb}^{*}$ $0.80\pm 0.08$ $1.080\pm 0.05$
@@
To clarify the impact of the present extended analysis, we provide a detailed comparison with previous QCD sum rule studies, as summarized in the tables 3, 4, 5, 6, 7 and 8. For the ground state, the masses obtained in this work are in good agreement with most existing QCD sum rule predictions. The only notable deviation appears in the $\Xi_{bc}^{*}$ channel, where our result differs from Ref. Zhang:2008rt , while consistency is observed with the remaining studies (Refs. Aliev:2012iv ; Wang:2010vn ). For the $1P$ excited states, our mass predictions are compatible with Ref. Wang:2010it . However, except for the $\Omega_{bb}^{*}$ channel, noticeable numerical differences are found when compared with Ref. Alrebdi:2022lat . A similar pattern is observed for the $2S$ states. Apart from the $\Omega_{bb}^{*}$ channel, our predicted masses show sizable deviations from those reported in Ref. Alrebdi:2022lat .
Regarding the pole residues, for the ground states our results differ from those reported in Refs. Aliev:2012iv ; Wang:2010vn ; Bagan:1992za in most channels, with agreement occurring only in a limited number of cases. For the $1P$ states, the extracted residues also exhibit differences compared to Refs. Wang:2010it ; Alrebdi:2022lat , except in a few specific channels. Furthermore, for the $2S$ states, our predicted residues differ numerically from Ref. Alrebdi:2022lat , with the exception of the $\Xi_{cc}^{*}$ and $\Omega_{cc}^{*}$ channels where reasonable agreement is observed.
The observed differences between our results and previous QCD sum rule analyses mainly originate from several methodological aspects. First, different choices of the auxiliary parameters, such as the Borel window and the continuum threshold $s_{0}$, can significantly affect the extracted observables. Second, the use of different interpolating currents for doubly heavy baryons modifies the structure of the correlation function and the relative pole continuum contributions. Third, the choice of renormalization scale $\mu$ plays an essential role in heavy–quark systems. In the present work, following standard QCD sum rule prescriptions, we vary the renormalization scale in the range $\mu=(2\text{--}4)~\mathrm{GeV}$. The final reported masses and residues correspond to the averaged values obtained within this window, and the associated scale dependence is included in the quoted uncertainties. Moreover, our operator product expansion is carried out consistently up to dimension ten condensates. In many previous QCD sum rule studies of these systems, the expansion was truncated at lower dimensions (typically dimension five or dimension six), which can partially explain the numerical differences observed here. In order to estimate the uncertainties associated with the factorization (vacuum saturation) approximation for higher-dimensional condensates (dimension $\geq 6$), we vary the parameter $\kappa$ in the range $\kappa=(1\text{--}5)$. The maximal variation induced by this procedure is incorporated into the total uncertainties of both the masses and the residues. Numerically, this effect contributes less than approximately $2\%$ to the total uncertainties.
@@
Accurate spectroscopic parameters, including masses and residues, are essential for investigating the interactions and decays of doubly heavy baryons. Theoretical calculations of these parameters also serve as crucial guides for experimental searches for yet-undiscovered baryon states predicted by quark models. This study focuses on precisely determining the spectroscopic parameters of doubly heavy baryons across the ground state ($1S$), first orbital excitation ($1P$), and the first radial excitation ($2S$). Using the QCD sum rule approach, a reliable and predictive technique based on the QCD Lagrangian, we include nonperturbative operators up to dimension ten to enhance accuracy. Following the determination of auxiliary parameters, we evaluate the mass spectra and residue for these specific baryon states, presenting them in table 2 and comparing our findings with results from other theoretical frameworks. The outcomes of this study are poised to aid experimental groups in their quest to discover the yet-unseen members of the doubly heavy baryon family. Additionally, these predictions offer a valuable basis for comparison with future theoretical work employing different nonperturbative methods. The importance of residue values, critical for computing any physical quantities related to baryon interactions and decays, cannot be overstated. Our results provide the necessary input data for comprehensive studies of the decay characteristics of the considered doubly heavy baryons in their ground and excited configurations. These investigations are essential for determining the decay widths and estimating the lifetimes of these baryons.